Let  . P(x): x2/2 = x


Find the following then identify their truth values.

  1. P (1)
  2. P (2)
  3. ꓯn, P(n)
  4. ⱻn, P(n)
1
Expert's answer
2022-04-06T16:38:47-0400

"1:\\\frac{1^2}{2}\ne 1\Rightarrow P\left( 1 \right) =0\\2:\\\frac{2^2}{2}=2\Rightarrow P\left( 2 \right) =1\\3:\\P\left( 1 \right) =0\Rightarrow \forall nP\left( n \right) =0\\4:\\P\left( 2 \right) =1\Rightarrow \exists nP\left( n \right) =1"